OpenStudy (anonymous):

FANN & MEDALL 1. Find the standard deviation of 4,8,5,12,3,9,5,2

OpenStudy (xapproachesinfinity):

did you find the mean yet?

OpenStudy (xapproachesinfinity):

\[\frac{4+8+5+12+3+9+5+2}{8}=?\]

OpenStudy (xapproachesinfinity):

that would be the mean

OpenStudy (anonymous):

the mean is 6

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

yes

OpenStudy (xapproachesinfinity):

now each number subtract the mean from it and square it

OpenStudy (xapproachesinfinity):

4-6=-2 (-2)^2=4 8-6=2 2^2=4 and continue

OpenStudy (xapproachesinfinity):

we are trying the find the variance

OpenStudy (anonymous):

okay... hold on

OpenStudy (xapproachesinfinity):

did you do it?

OpenStudy (anonymous):

5-6=-1 (-1)^2=-1 12-6=6 (6)^2=36 3-6=-3 (-3)^2=-9 9-6=3 (3)^2=9 5-6=-1 (-1)^2=-1 2-6=-4 (-4)^2=-16

OpenStudy (anonymous):

right? @xapproachesinfinity

OpenStudy (xapproachesinfinity):

(-1)^2=1 not -1 (-3)^2=9 not -9 same for the rest

OpenStudy (xapproachesinfinity):

when you square a number negative vanishes

OpenStudy (xapproachesinfinity):

so we got 4,4,1,36,9,9,1,16 now we need to find the mean for this numbers

OpenStudy (xapproachesinfinity):

see how i found the mean for the first step

OpenStudy (anonymous):

yes... do i find it for those numbers now?

OpenStudy (xapproachesinfinity):

yes

OpenStudy (anonymous):

okay.. hold on

OpenStudy (anonymous):

10?

OpenStudy (xapproachesinfinity):

yes

OpenStudy (xapproachesinfinity):

now take the square root of that

OpenStudy (anonymous):

3.16?

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (anonymous):

is that gonna be the answer?

OpenStudy (xapproachesinfinity):

yes

OpenStudy (anonymous):

thxx so much!!! :)

OpenStudy (xapproachesinfinity):

no problem